Austin wants to glue string diagonally across the cover of his rectangular box which is 30 cm in lenghth and 16 cm in width. How many cm of string will Austin need for both diagonals of the box cover?
Choices are: A) 34 cm B) 14 cm C) 68 cm D) 28cm
I tried h+h+w+w and not right
h+h+w not right
Please help me with the formula

Answer by edjones(8007)

2sqrt(30^2+16^2)=68 a^2+b^2=c^2